A medical statistician wants to estimate the average weight loss of people who are on a new diet plan. In a preliminary study, he guesses that the standard deviation of the population of weight losses is about 10 pounds. How large a sample should he take to estimate the mean weight loss to within 2 pounds, with a 90% confidence?



1
Expert's answer
2021-07-11T09:49:59-0400

The critical value for α=0.1\alpha=0.1 is zc=z1α/2=1.6449z_c=z_{1-\alpha/2}=1.6449


SE=zcσnSE=z_c\cdot\dfrac{\sigma}{\sqrt{n}}

Given σ=10,SE=2\sigma=10, SE=2


1.644910n21.6449\cdot\dfrac{10}{\sqrt{n}}\leq2

n(1.6449102)2n\geq(1.6449\cdot\dfrac{10}{2})^2

n68n\geq68

The total sample should be at least 68.